OpenStudy (anonymous):

Complete the square. Solve for x: (x - 9)2 = 1.

OpenStudy (paxpolaris):

is that :\[\left( x-9 \right)^{2}=1\]

OpenStudy (cathyangs):

x-9 is either equal to -1 or 1. write two equations, and solve both.

OpenStudy (anonymous):

Yes Pax

OpenStudy (paxpolaris):

cathy is right... reversing the squares on both sides you would get: \[x-9=\pm1\]\[x=\pm1+9\] So, \[x= 1+ 9 = 10\] or \[x= -1+ 9 = 8\] x= 8,10